Two loudspeakers emit sound waves along the x-axis. The sound has maximum intensity when the speakers are 42 cm apart. The sound intensity decreases as the distance between the speakers is increased, reaching zero at a separation of 49 cm .

What is the wavelength of the sound?

If the distance between the speakers continues to increase, at what separation will the sound intensity again be a maximum?
Answer #1

Here,

Highest intensity = 42 cm
Lowest intensity = 49 cm

As,49-42 = 7 cm. But that's 7 cm for half a wave so *2 = 14 cm = wavelength.

From 49 cm to another point of high intensity = 49+ 14 = 63 cm
Answer #2

To do part b correctly you take the distance from 49-42 = 7cm and add it to the largest distance of 49cm. 


Therefore we have: 49cm+7cm = 56cm

answered by: anonymous
Answer #3

image.png

answered by: Allen